# H2R: A Human-to-Robot Data Augmentation for Robot Pre-training from Videos

### v3

@@ 4 Simulation Experiment
Simulation Benchmark. For each pre-training method, we evaluate the performance of pre-trained encoders in imitation learning. Specifically, we select a total of seven simulation tasks in different environments, which are from Robomimic [mandlekar2021matters], RLbench [james2019rlbench], and Diffusion Policy [chi2023diffusionpolicy]. In particular, for Robomimic, we train the policies using the behavior cloning (BC) and evaluate them on tasks such as MoveCan, Square, and Lift, where the robot performs actions such as moving or lifting objects. For RLBench, we train the policies with Diffusion Policy and evaluate them on three manipulation tasks: Close Box, Close Laptop Lid and Toilet Seat Down. We use the PushT task in the Diffusion Policy evaluation framework, which evaluates a robot’s ability to push an object to a target location. We also use three MetaWorld tasks: Assembly, ButtonPress and Hammer from VC-1 Cortex benchmark for evaluation. An overview of these simulation tasks is shown in Appendix.
Table 2: Simulation Benchmark Results. Success rates (%) across imitation learning tasks in Robomimic, PushT, RLBench and CortexBench, evaluated with MAE and R3M encoders before and after applying H2R. Blue represents an increase in task success rate, while Red represents a decrease. All subsequent tables follow the same rule.
Robomimic Diffusion Policy RLBench CortexBench MoveCan Square Lift Average PushT CloseBox CloseLaptopLid ToiletSeatDown Average Assembly ButtonPress Hammer Average MAE (SSv2) 54.0 25.5 94.5 58.0 59.2 0.0 10.0 0.0 3.3 84.0 80.0 96.0 86.7 MAE (H2R-SSv2) 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) 5.0 (+5.0%) 15.0 (+5.0%) 20.0 (+20.0%) 13.3 (+10.0%) 88.0(+4.0%) 88.0(+8.0%) 100.0 (+4.0%) 92.0(+5.3%) R3M (SSv2) 59.5 20.5 85 55 15 0 20 10 10 76.0 56.0 88 73.3 R3M (H2R-SSv2) 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%) 5.0 (+5.0%) 20.0 (0.0%) 20.0 (+10.0%) 15.0 (+5.0%) 68.0 (-8.0%) 60.0 (+4.0%) 96.0 (+8.0%) 74.7 (+1.3%)
4.1 Simulation Results
Table 2 shows encoders trained on augmented data H2R-SSv2 significantly outperform those on original SSv2 across all simulation tasks. For Robomimic tasks, MAE achieves a 10.2% average success rate improvement and R3M 6.3%, with MAE seeing a substantial 25.5% gain on MoveCan. On PushT, H2R boosts MAE by 5.3% and R3M by 7.0%. In RLBench, MAE improves its performance by 10.0% on average and R3M by 5.0%. In CortexBench, H2R improves MAE by 5.3% and R3M by 1.3%. These results confirm that the H2R-augmented dataset delivers superior visual representations for imitation learning to those learned from large-scale human video datasets like SSv2, as H2R  bridges the visual gap between human hands and robotic arms to enhance vision encoder pre-training performance.
Table 3: Results on Ego4D Dataset. Success rates (%) across imitation learning tasks in the PushT and RLBench environments.
Diffusion Policy RLBench PushT CloseBox CloseLaptopLid ToiletSeatDown Average MAE (Ego4D) 51.3 0 0 5 1.7 MAE (H2R-Ego4D) 53.5 (+2.2%) 10 (+10%) 5 (+5.0%) 0 (-5.0%) 5 (+3.3%) R3M (Ego4D) 13.6 10 5 5 6.7 R3M (H2R-Ego4D) 13.6 (+0.0%) 15 (+5.0%) 5 (0.0%) 15.0 (+10.0%) 11.7 (+10.0%)
In addition to SSv2 dataset, we also perform the same experiments on the PushT task and the tasks in RLbench benchmark using the Ego4D dataset. The experimental results in the simulator are shown in Table 3, which indicates that H2R remains highly effective when applied to the Ego4D dataset.
4.2 Pretraining Model on Robotic Datasets
Table 4: Comparison with model pretrained on robotic datasets. All visual encoder weights are initialized using the R3M pretraining method, with the “R3M-H2R” variant pretrained on the H2R-augmented SSv2 dataset.
Robomimic MoveCan Square Lift Average R3M 59.5 20.5 85.0 55.0 R3M-H2R 61.5 37.5 85.0 61.3 R3M-DROID 54.0 22.0 96.0 56.7
To further evaluate how H2R compares to models pretrained directly on robotic datasets, we refer to a recent study [jiang2024robotspretrainrobotsmanipulationcentric], which conducted a systematic comparison between visual representations pretrained on large-scale egocentric human video datasets and those trained on robotic datasets such as DROID [khazatsky2024droidlargescaleinthewildrobot]. The experimental results, summarized in Table 4, show that models pretrained on robotic datasets often provide limited or even degraded performance on downstream manipulation Robomimic benchmarks. Moreover, despite being smaller and simulated rather than real-robot data, the H2R dataset outperforms the DROID dataset in pre-training effectiveness, indicating its strong potential for robotic model pre-training and reflecting its high data quality.
4.3 Effects of Demonstration Density
The relatively low success rates observed in Table 2 are primarily due to our use of keypoint frames in RLBench tasks. This approach, widely adopted in the RLBench benchmark [goyal2023rvt, jia2024lift3dfoundationpolicylifting, shridhar2023peract], accelerates both training and evaluation by reducing redundant frames. However, selecting only keyframes inevitably decreases the density of demonstration data. Since the diffusion policy employed in our imitation learning framework benefits from dense and temporally continuous data due to its long observation horizon, sparse sampling may limit its performance.
To further investigate demonstration density impact and validate H2R’s robustness, we retrain the diffusion policy with a dense RLBench dataset under identical settings. Table 5 shows higher data density significantly boosts overall success rate (e.g., MAE average from 3.3% to 45.0%). Critically, H2R consistently enhances performance across sparse and dense data regimes, confirming its general effectiveness regardless of demonstration density.
Table 5: Effects of Demonstration Density. Success rates (%) across imitation learning tasks in the RLBench environments with sparse data and dense data.
Data Model RLBench CloseBox CloseLaptopLid ToiletSeatDown Average Sparse data MAE 0.0 10.0 0.0 3.3 MAE+H2R 5.0 (+5.0%) 15.0 (+5.0%) 20.0 (+20.0%) 13.3 (+10.0%) Dense data MAE 50.0 40.0 45.0 45.0 MAE+H2R 60.0 (+10.0%) 45.0 (+5.0%) 60.0 (+15.0%) 55.0 (+10.0%)
5 Real World Experiment
Table 6: Real-world Task Results. We report the success rate (%) over real-world tasks for MAE and R3M.
Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) DP Gripper-PickCube 45 65 (+20%) 40 50 (+10%) ACT Gripper-PickCube 25 30 (+5%) 25 30 (+5%) Gripper-Stack 50 55 (+5%) 55 70 (+15%) Gripper-Stack 20 35 (+15%) 20 40 (+20%) Gripper-CloseBox 55 50 (-5%) 45 65 (+20%) Gripper-CloseBox 35 40 (+5%) 40 50 (+10%) Average 50 56.7 (+6.7%) 46.7 61.7 (+15%) Average 26.7 35 (+8.3%) 28.3 40 (+11.7%) Leaphand-GraspChicken 40 55 (+15%) 10 35 (+25%) Leaphand-GraspChicken 45 50 (+5%) 10 35 (+25%) Leaphand-StandCup 35 60 (+25%) 20 50 (+30%) Leaphand-StandCup 25 50 (+25%) 20 60 (+40%) Leaphand-OpenBox 45 65 (+20%) 40 45 (+5%) Leaphand-OpenBox 30 40 (+10%) 15 20 (+5%) Average 40 60 (+20%) 23.3 43.3 (+20.0%) Average 33.3 46.7 (+13.3%) 15 38.3 (+23.3%) Franka-PlaceToy 10 15 (+5%) 10 25 (+15%) Franka-PlaceToy 25 40 (+15%) 20 30 (+10%) Franka-SweepRubbish 15 15 (0%) 20 20 (0%) Franka-SweepRubbish 20 35 (+15%) 20 20 (0%) Franka-WeightSauce 40 45 (+5%) 20 30 (+10%) Franka-WeightSauce 30 35 (+25%) 20 25 (+5%) Average 21.7 25 (+3.3%) 17.7 25 (+7.3%) Average 25 36.7 (+16.7%) 20 25 (+5%)
Table 7: Real-world Success Rates with H2R. We report the task success rates (%) using two robot embodiments (UR5, Franka) for data augmentation during pre-training. All models are evaluated on downstream tasks using the Leaphand end-effector, under MAE and R3M encoders with DP and ACT policies. Despite embodiment mismatch, H2R provides consistent gains over the baseline without augmentation.
Policy Task MAE (SSv2) MAE (H2R-UR5-SSv2) MAE (H2R-Franka-SSv2) R3M(SSv2) R3M (H2R-UR5-SSv2) R3M (H2R-Franka-SSv2) DP Leaphand-GraspChicken 40 55 (+15%) 35 (-5%) 10 35 (+25%) 20 (+10%) Leaphand-StandCup 35 60 (+25%) 50 (+15%) 20 50 (+30%) 30 (+10%) Leaphand-OpenBox 45 65 (+20%) 45 40 45 (+5%) 45 (+5%) Average 40 60 (+20%) 43.3 (+3.3%) 21.7 43.3 (+21.7%) 31.7 (+10%) ACT Leaphand-GraspChicken 45 50 (+5%) 50 (+5%) 10 35 (+25%) 40 (+30%) Leaphand-StandCup 25 50 (+25%) 50 (+25%) 20 60 (+40%) 25 (+5%) Leaphand-OpenBox 30 40 (+10%) 30 15 20 (+5%) 5 (-10%) Average 33.3 46.7 (+13.3%) 43.3 (+10%) 15 38.3 (+23.3%) 23.3 (+8.3%)
Robot Setups. We evaluate the effectiveness of H2R across three real-world manipulation setups: (i) a UR5 arm equipped with a Robotiq Gripper [RobotiqAdaptiveGrippers], (ii) a UR5 arm equipped with a Leaphand end effector [shaw2023leaphandlowcostefficient], and (iii) a dual-arm Franka Emika system using parallel Grippers. Realsense [IntelRealSenseL515] is mounted on the side of the robotic arm, which provides a similar viewpoint to the ego-centric human video data used in the pre-trained visual model. We provide the visualization of the robot setup in real-world scenarios in the supplementary materials.
Figure 4: Visualization of Real-world Manipulation Tasks. The left columns show Gripper tasks and the right columns show Leaphand tasks. Each task is illustrated with three frames, demonstrating the progression from the initial state to the completion of the manipulation.
Real-world Tasks. We set up nine tasks for gripper manipulation, dexterous manipulation and dual-arm manipulation (Figure 4). For the Gripper tasks, we design: (1) Gripper-PickCube, where the Gripper picks up a cube and places it into a bowl; (2) Gripper-Stack, where a blue cube is stacked atop a yellow cube; and (3) Gripper-CloseBox, where a cube is retrieved from a box, placed into a bowl, and the box lid is subsequently closed. For the Leaphand tasks, we design: (4) Leaphand-GraspChicken, where a toy chicken is grasped and placed into a bowl; (5) Leaphand-StandCup, where a fallen cup is stood upright on the table; and (6) Leaphand-OpenBox, where an articulated box lid is opened. For the dual-arm manipulation tasks, we design: (7) Dualarm-PlaceToy, where the left arm takes a toy from the pink box and the right arm moves it to the blue box; (8) Dualarm-SweepRubbish, where the left arm sweeps the trash into the dustpan held by the right arm; and (9) Dualarm-WeightSauce, where the left arm places the cup on the scale and the right arm picks up the bottle for a pouring motion. We adopt varying numbers of demonstrations, episode lengths per demonstration, and maximum action steps during evaluation for each task, as in Table 8.
Table 8: Task-specific Data Collection and Evaluation Settings.
Task Episode Length Num Demos Max Steps Gripper-PickCube 45 30 80 Gripper-Stack 43 30 80 Gripper-CloseBox 68 30 120 Leaphand-GraspChicken 150 50 500 Leaphand-StandCup 150 50 500 Leaphand-OpenBox 200 50 800 Dualarm-PlaceToy 1000 300 1500 Dualarm-SweepRubbish 1000 300 1500 Dualarm-WeightSauce 1000 300 1500
Table 9: Ablation Study. Ablation by removing robot overlay (w/o Overlay) and camera-hand retargeting (w/o Retarget).
Policy Task H2R H2R w/o Overlay H2R w/o Retarget DP Leaphand-GraspChicken 55 30 (-25%) 30 (-25%) Leaphand-StandCup 60 40 (-20%) 55 (-5%) Leaphand-OpenBox 65 20 (-45%) 45 (-20%) Average 60 30 (-30%) 43.3 (-16.7%) ACT Leaphand-GraspChicken 50 25 (-25%) 45 (-5%) Leaphand-StandCup 50 35 (-15%) 30 (-20%) Leaphand-OpenBox 40 25 (-15%) 15 (-25%) Average 46.7 28.3 (-18.3%) 30 (-16.7%)
@@ 5 Real World Experiment
Policy Training. For policy training, we select the Diffusion Policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act] as policy frameworks. We apply the pre-trained MAE and R3M visual encoders to downstream policy learning, following the same pretraining configuration described in Section 4.
@@
Evaluation. To ensure that the initial conditions are consistent throughout the evaluation, we randomly place the target objects within a predefined area following a uniform distribution during expert demonstrations. During the evaluation, each real-world task is rolled out 20 times. We report the success rates of these tasks to assess model quality.
5.1 Real-world Results
In the real-world tasks, we employ diffusion policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act] to drive robots in performing manipulation tasks, where the visual encoders are pre-trained using MAE and R3M. To evaluate the effectiveness and generalization of H2R, we first conduct experiments to validate its performance enhancement when the pre-training and policy training embodiments are the same. We then perform cross-embodiment experiments to investigate the transfer effects of H2R when the pre-training and policy training embodiments differ. We finally apply H2R to fine-tune a pre-trained vision backbone in a Vision-Language-Action model, showcasing its promising potential in this advanced setting.
For identical embodiment experiment, we use UR5 or Franka robot for augmentation during pre-training, which matches the embodiment used in policy training and evaluation. As shown in Table 6, H2R provides consistent improvements across both MAE and R3M encoders for both DP and ACT policies. For Leaphand tasks under DP, H2R leads to an average improvement of 20% (MAE) and 21.7% (R3M). For Gripper tasks, improvements reach 6.7% (MAE) and 15% (R3M). For Dual-arm Franka tasks, the improvements reach 3.3% (MAE) and 7.3% (R3M). Similar trends are observed in ACT, with Leaphand performance increasing by 13.3% (MAE) and 23.3% (R3M), Gripper tasks improving by 8.3%(MAE) and 11.7%(R3M), and Dual-arm Franka tasks improving by 16.7%(MAE) and 5%(R3M).
To evaluate cross-embodiment generalization, we use Franka-augmented the SSv2 dataset with H2R pipeline. Downstream policy training and evaluation are conducted with UR5 Leaphand. As shown in Table 7, H2R still outperforms raw Ego4D despite the embodiment mismatch. Under DP, MAE improves from 40% to 43.3%, and R3M from 21.7% to 31.7%. Under ACT, MAE increases from 33.3% to 43.3%, and R3M from 15% to 23.3%.
5.2 Performance on Other Datasets
@@
Table 10: Real-world Success Rates with Ego4D Pre-training. We report success rates (%) on real-world Leaphand tasks using visual encoders pre-trained on the Ego4D dataset. All downstream policies are trained using ACT.
Task MAE (Ego4D) MAE (H2R-Ego4D) R3M (Ego4D) R3M (H2R-Ego4D) Leaphand-GraspChicken 25 35 (+10%) 15 20 (+5%) Leaphand-StandCup 25 50 (+25%) 20 25 (+5%) Leaphand-OpenBox 35 45 (+10%) 30 40 (+10%) Average 28.3 43.3 (+15%) 21.7 28.3 (+6.7%)
We compare encoders pre-trained on the original Ego4D dataset and on the H2R-augmented dataset, and apply them to real-world Leaphand manipulation tasks using the ACT [zhao2023act] policy. As shown in Table 10, models trained with H2R data consistently outperform those using raw Ego4D across all three Leaphand manipulation tasks, achieving an average improvement of 15% for MAE and 6.7% for R3M. These results complement our findings on SSv2 and reinforce that robot-centric augmentation improves downstream real-world performance.
5.3 Potential for VLA Model Pretraining
Figure 5: Dual-UR5 tasks. The left columns shows PlaceBowl task and the right columns shows StackBlocks task.

### v4

@@ IV-A Experimental Setup
Simulation Benchmark. For each pre-training method, we evaluate the performance of pre-trained encoders in imitation learning. Specifically, we select a total of 10 simulation tasks in different environments, which are from Robomimic [36], RLBench [24], PushT [7] and CortexBench [35]. In particular, for Robomimic, we train the policies using the behavior cloning (BC) and evaluate them on tasks such as MoveCan, Square, and Lift, where the robot performs actions such as moving or lifting objects. For RLBench, we train the policies with Diffusion Policy and evaluate them on three manipulation tasks: Close Box, Close Laptop Lid and Toilet Seat Down. We use the PushT task in the Diffusion Policy evaluation framework, which evaluates a robot’s ability to push an object to a target location. We also use three MetaWorld tasks: Assembly, ButtonPress and Hammer from VC-1 Cortex benchmark for evaluation.
Real-world Setup.
TABLE II: Real-world Manipulation Tasks. Overview of task definitions grouped by robot embodiment.
Embodiment Task Name Task Description UR5-Gripper PickCube The robot grasps a cube and places it into a bowl. Stack The robot stacks a blue cube on top of a yellow cube. CloseBox The robot retrieves a cube from a box, places it into a bowl, and closes the box lid. UR5-Leaphand GraspChicken The robot grasps a toy chicken and places it into a bowl. StandCup The robot grasps a fallen cup and places it upright on the table. OpenBox The robot opens an articulated box lid using the dexterous hand. Dual-arm Franka PlaceToy The left arm grasps a toy from a pink box, and the right arm transfers it into a blue box. SweepRubbish The left arm sweeps trash into a dustpan held by the right arm. WeighSauce The left arm places a cup onto a scale, while the right arm performs a pouring motion with a bottle. Dual-arm UR5e PlaceBowl The robot places a green bowl on top of a blue bowl. StackBlocks The robot stacks blocks in the order of red, yellow, and blue to form a single tower.
We evaluate the effectiveness of H2R across four real-world manipulation setups: (i) a UR5 arm equipped with a Robotiq Gripper [46], (ii) a UR5 arm equipped with a Leaphand end effector [49], (iii) a dual-arm Franka Emika system with parallel grippers, and (iv) a dual-arm UR5e platform configured as a human-like bimanual manipulation setup. Across these platforms, we evaluate a diverse set of real-world manipulation tasks spanning gripper-based, dexterous, and bimanual manipulation; detailed task definitions are summarized in Table II.
Demonstration data are collected through human teleoperation, with the collection interface adapted to the manipulation setting. Specifically, gripper-based tasks are collected using keyboard-based teleoperation, Leaphand tasks using vision-guided teleoperation following the same setup as in CordViP [16], and dual-arm tasks using homogeneous-arm bilateral teleoperation [54]. The number of demonstrations varies across setups and task categories, ranging from 30 to 300 episodes depending on task complexity, as summarized in Table I.
@@
For policy training, we adopt Diffusion Policy (DP) [7] and ACT [57] as policy frameworks. We apply the pre-trained MAE and R3M visual encoders to downstream policy learning.
@@
During evaluation, target objects are randomly initialized within predefined regions following a uniform distribution consistent with expert demonstrations. Each task is executed for 20 rollouts, and success rates are reported as the primary evaluation metric.
IV-B Simulation Results
Performance on Simulation Benchmarks.
TABLE III: Simulation Benchmark Results. Success rates ($\uparrow$) across diverse imitation learning suites. Bold indicates the best performance within each group, and green/red denotes the performance gain/drop after applying H2R. All subsequent tables follow the same rule.
Method Robomimic PushT RLBench CortexBench MoveCan Square Lift Avg. PushT CloseBox CloseLaptopLid ToiletSeatDown Avg. Assembly ButtonPress Hammer Avg. MAE (SSv2) 54.0 25.5 94.5 58.0 59.2 0.0 10.0 0.0 3.3 84.0 80.0 96.0 86.7 MAE (H2R) 79.5 29.5 95.5 68.2 64.5 5.0 15.0 20.0 13.3 88.0 88.0 100.0 92.0 Gain ($\Delta$) +25.5 +4.0 +1.0 +10.2 +5.3 +5.0 +5.0 +20.0 +10.0 +4.0 +8.0 +4.0 +5.3 R3M (SSv2) 59.5 20.5 85.0 55.0 15.0 0.0 20.0 10.0 10.0 76.0 56.0 88.0 73.3 R3M (H2R) 61.5 37.5 85.0 61.3 22.0 5.0 20.0 20.0 15.0 68.0 60.0 96.0 74.7 Gain ($\Delta$) +2.0 +17.0 0.0 +6.3 +7.0 +5.0 0.0 +10.0 +5.0 -8.0 +4.0 +8.0 +1.3
Table III shows that visual encoders pre-trained with H2R consistently outperform those trained on the original SSv2 dataset across all evaluated simulation benchmarks for both MAE and R3M visual encoders. For Robomimic tasks, H2R leads to clear improvements for both encoders. MAE achieves a 10.2% increase in average success rate, while R3M improves by 6.3%. Notably, MAE observes a substantial 25.5% gain on the MoveCan task. Similar trends are observed on PushT, RLBench, and CortexBench, with consistent performance gains across the evaluated tasks for both MAE and R3M. These results demonstrate that H2R augmentation consistently enhances the effectiveness of visual representations learned from large-scale human video datasets for downstream imitation learning across diverse simulation environments.
Generalization Across Pre-training Datasets.
TABLE IV: Simulation Results with Ego4D Pre-training. Success rates ($\uparrow$) on PushT and RLBench when pre-training MAE/R3M on an Ego4D subset, comparing original vs. H2R-augmented pre-training.
Method PushT RLBench PushT CloseBox CloseLaptopLid ToiletSeatDown Avg. MAE (Ego4D) 51.3 0.0 0.0 5.0 1.7 MAE (H2R) 53.5 10.0 5.0 0.0 5.0 Gain ($\Delta$) +2.2 +10.0 +5.0 -5.0 +3.3 R3M (Ego4D) 13.6 10.0 5.0 5.0 6.7 R3M (H2R) 13.6 15.0 5.0 15.0 11.7 Gain ($\Delta$) 0.0 +5.0 0.0 +10.0 +5.0
We evaluate H2R using Ego4D to assess its generalization beyond SSv2. Specifically, we pre-train both MAE and R3M on an Ego4D subset of comparable scale to SSv2, following the same training protocol as in the previous experiments. Table IV reports the resulting imitation learning performance on the PushT task and RLBench benchmarks. On PushT, H2R leads to improvement for MAE (+2.2%) while preserving the same performance for R3M. On RLBench, H2R improves the average success rate for both encoders, with MAE increasing from 1.7% to 5.0% and R3M from 6.7% to 11.7%.
These results indicate that the benefits of H2R are not specific to SSv2, and can be consistently transferred to visual representations pre-trained on Ego4D, supporting the general applicability of H2R across different sources of large-scale human video data.
Pretraining on Robotic Datasets. We further compare H2R with visual representations pretrained directly on robotic datasets by including models pretrained on the DROID dataset [29], following the setting reported in the recent study [26]. All models use the R3M framework for pre-training, where R3M is pretrained on SSv2, R3M-DROID is pretrained on the DROID robotic dataset, and R3M-H2R is pretrained on the H2R-augmented SSv2 dataset. We evaluate all representations on the same Robomimic benchmarks. As shown in Table V, both robotic data pre-training and H2R augmentation lead to performance improvements over SSv2-only pre-training. However, R3M-H2R achieves the strongest overall results. Despite DROID being collected from real-world robotic executions, H2R pre-training achieves stronger downstream manipulation performance, suggesting that human-to-robot visual augmentation offers an effective alternative to direct robotic data pretraining.
TABLE V: Comparison to Robotic-data Pre-training. Robomimic success rates ($\uparrow$) for R3M pre-training on SSv2, robotic dataset (DROID), and H2R-augmented SSv2, highlighting H2R as an alternative to direct robotic-data pre-training.
Method MoveCan Square Lift Avg. R3M 59.5 20.5 85.0 55.0 R3M-DROID 54.0 22.0 96.0 56.7 R3M-H2R 61.5 37.5 85.0 61.3
@@ IV-B Simulation Results
TABLE VI: Real-world Task Results. Success rates ($\uparrow$) on gripper, dexterous (Leaphand), and bimanual (Franka) tasks using DP and ACT policies with MAE/R3M visual encoders. Rows denote pre-training variants, and columns denote tasks.
(a) DP Policy.
Policy Method Tasks Gripper Leaphand Franka PickCube Stack CloseBox Avg. GraspChicken StandCup OpenBox Avg. PlaceToy SweepRubbish WeighSauce Avg. DP MAE (SSv2) 45.0 50.0 55.0 50.0 40.0 35.0 45.0 40.0 10.0 15.0 40.0 21.7 MAE (H2R) 65.0 55.0 50.0 56.7 55.0 60.0 65.0 60.0 15.0 15.0 45.0 25.0 Gain ($\Delta$) +20.0 +5.0 -5.0 +6.7 +15.0 +25.0 +20.0 +20.0 +5.0 0.0 +5.0 +3.3 R3M (SSv2) 40.0 55.0 45.0 46.7 10.0 20.0 40.0 23.3 10.0 20.0 20.0 17.7 R3M (H2R) 50.0 70.0 65.0 61.7 35.0 50.0 45.0 43.3 25.0 20.0 30.0 25.0 Gain ($\Delta$) +10.0 +15.0 +20.0 +15.0 +25.0 +30.0 +5.0 +20.0 +15.0 0.0 +10.0 +7.3
(b) ACT Policy.
Policy Method Tasks Gripper Leaphand Franka PickCube Stack CloseBox Avg. GraspChicken StandCup OpenBox Avg. PlaceToy SweepRubbish WeighSauce Avg. ACT MAE (SSv2) 25.0 20.0 35.0 26.7 45.0 25.0 30.0 33.3 25.0 20.0 30.0 25.0 MAE (H2R) 30.0 35.0 40.0 35.0 50.0 50.0 40.0 46.7 40.0 35.0 35.0 36.7 Gain ($\Delta$) +5.0 +15.0 +5.0 +8.3 +5.0 +25.0 +10.0 +13.3 +15.0 +15.0 +5.0 +11.7 R3M(SSv2) 25.0 20.0 40.0 28.3 10.0 20.0 15.0 15.0 20.0 20.0 20.0 20.0 R3M (H2R) 30.0 40.0 50.0 40.0 35.0 60.0 20.0 38.3 30.0 20.0 25.0 25.0 Gain ($\Delta$) +5.0 +20.0 +10.0 +11.7 +25.0 +40.0 +5.0 +23.3 +10.0 0.0 +5.0 +5.0
TABLE VII: Effects of Demonstration Density. Success rates ($\uparrow$) under sparse vs. dense RLBench demonstrations, comparing MAE encoders pre-trained with and without H2R.
Data Model CloseBox CloseLaptopLid ToiletSeatDown Avg. Sparse data MAE 0.0 10.0 0.0 3.3 MAE (H2R) 5.0 15.0 20.0 13.3 Gain ($\Delta$) +5.0 +5.0 +20.0 +10.0 Dense data MAE 50.0 40.0 45.0 45.0 MAE (H2R) 60.0 45.0 60.0 55.0 Gain ($\Delta$) +10.0 +5.0 +15.0 +10.0
Effects of Demonstration Density. We observe relatively low success rates on RLBench under sparse, keypoint-based supervision, a setting commonly adopted in prior RLBench works to improve training efficiency [18, 25, 50]. This observation motivates an additional study to examine whether the effectiveness of H2R depends on demonstration density, by comparing sparse and dense RLBench training under identical settings.
As shown in Table VII, increasing demonstration density substantially improves overall task performance (e.g., MAE average success rate increases from 3.3% to 45.0%). More importantly, H2R consistently yields performance gains in both sparse and dense regimes. This result indicates that H2R is not tailored to a specific data density, but instead provides robust visual representations that remain effective across different demonstration settings.
IV-C Real-world Results
Performance on Real-world Manipulation Tasks. In real-world experiments, we evaluate H2R on three categories of manipulation tasks-gripper, dexterous, and bimanual—using corresponding robotic platforms, including a UR5 equipped with a Robotiq gripper, a UR5 equipped with a Leaphand end effector, and a dual-arm Franka system. We adopt Diffusion Policy (DP) [7] and ACT [57] as policy frameworks, with visual encoders pre-trained using MAE and R3M. In this setting, the robot embodiment used for H2R augmentation during pre-training is identical to that used in downstream policy training and evaluation.
As shown in Table VI, H2R consistently improves real-world success rates across all task categories, encoders, and policies. Notably, the most pronounced gains are observed on Leaphand tasks, where H2R improves average performance by 20.0% (MAE, R3M) under DP, and by 13.3% (MAE) and 23.3% (R3M) under ACT. Gripper-based and dual-arm Franka tasks also benefit from H2R, with consistent improvements across both MAE and R3M. These results demonstrate that H2R effectively enhances real-world manipulation performance.
TABLE VIII: Real-world results with Ego4D pre-training. Success rates ($\uparrow$) on UR5-Leaphand tasks using ACT, comparing Ego4D pre-training with vs. without H2R augmentation.
Method GraspChicken StandCup OpenBox Avg. MAE (Ego4D) 25.0 25.0 35.0 28.3 MAE (H2R) 35.0 50.0 45.0 43.3 Gain ($\Delta$) +10.0 +25.0 +10.0 +15.0 R3M (Ego4D) 15.0 20.0 30.0 21.7 R3M (H2R) 20.0 25.0 40.0 28.3 Gain ($\Delta$) +5.0 +5.0 +10.0 +6.7
TABLE IX: Cross-Embodiment Real-world Results. UR5-Leaphand task success rates ($\uparrow$) when the robot embodiment used for H2R augmentation during pre-training differs from the downstream embodiment (UR5-based vs. Franka-based augmentation).
(a) DP Policy.
Policy Method GraspChicken StandCup OpenBox Avg. DP MAE (SSv2) 40.0 35.0 45.0 40.0 MAE (H2R-UR5) 55.0 60.0 65.0 60.0 Gain ($\Delta$) +15.0 +25.0 +20.0 +20.0 MAE (H2R-Franka) 35.0 50.0 45.0 43.3 Gain ($\Delta$) -5.0 +15.0 0.0 +3.3 R3M (SSv2) 10.0 20.0 40.0 23.3 R3M(H2R-UR5) 35.0 50.0 45.0 43.3 Gain ($\Delta$) +25.0 +30.0 +5.0 +20.0 R3M(H2R-Franka) 20.0 30.0 45.0 31.7 Gain ($\Delta$) +10.0 +10.0 +5.0 +10.0
(b) ACT Policy.
Policy Method GraspChicken StandCup OpenBox Avg. ACT MAE (SSv2) 45.0 25.0 30.0 33.3 MAE (H2R-UR5) 50.0 50.0 40.0 46.7 Gain ($\Delta$) +5.0 +25.0 +10.0 +13.3 MAE (H2R-Franka) 50.0 50.0 30.0 43.3 Gain ($\Delta$) +5.0 +25.0 0.0 +10.0 R3M (SSv2) 10.0 20.0 15.0 15.0 R3M (H2R-UR5) 35.0 60.0 20.0 38.3 Gain ($\Delta$) +25.0 +40.0 +5.0 +23.3 R3M (H2R-Franka) 40.0 25.0 5.0 23.3 Gain ($\Delta$) +30.0 +5.0 -10.0 +8.3
Generalization Across Pre-training Datasets. To examine whether the effectiveness of H2R observed in simulation also extends to real-world settings, we further evaluate visual encoders pre-trained on Ego4D in real-world Leaphand manipulation tasks. Following the same pre-training protocol as in the simulation experiments, both MAE and R3M are trained on an Ego4D subset of comparable scale, with and without H2R augmentation, and deployed using the ACT policy.
Table VIII reports the success rates across three real-world Leaphand tasks. For MAE, applying H2R consistently improves performance on all tasks, increasing the average success rate from 28.3% to 43.3%. For R3M, H2R also yields consistent gains, with the average success rate improving from 21.7% to 28.3%.
These results are consistent with the trends observed in simulation and indicate that the benefits of H2R extend to real-world manipulation when visual encoders are pre-trained on Ego4D.
Cross-embodiment Generalization. We further evaluate the cross-embodiment generalization of H2R by applying H2R-augmented pre-training using a robot embodiment that differs from the one used in downstream policy training and evaluation. Specifically, we consider two augmentation settings during pre-training: UR5-based H2R and Franka-based H2R, while all downstream policies are trained and evaluated on UR5-Leaphand tasks.
As shown in Table IX, H2R continues to improve performance under embodiment mismatch for both MAE and R3M encoders, as well as for both DP and ACT policies. Using UR5-based H2R augmentation yields the strongest gains, with average success rates increasing by 20.0% (MAE) and 21.7% (R3M) under DP, and by 13.3% (MAE) and 23.3% (R3M) under ACT. When H2R is applied using Franka-based augmentation, performance improvements remain consistent but are generally smaller in magnitude.
These results suggest that the effectiveness of H2R is not strictly limited to cases where the pre-training and downstream embodiments are identical. Within the evaluated settings, H2R remains beneficial even when the augmentation embodiment differs from that used in policy learning, indicating a certain degree of robustness to embodiment variation.